# Physics motivated models of pulsar X-ray hotspots: off-center dipole configurations

### v1

@@ 4 Real case study: PSR J0030+0451 and PSR J0437–4715 hotspot map construction
In this case study, we focus on two sources: PSR J0030+0451 (J0030) and PSR J0437–4715 (J0437). PSR J0030+0451 was the first millisecond pulsar that NICER source for which detailed pulse profile modeling was performed. The original report on the hotspot configuration of this star, by Riley et al. (2019) and Miller et al. (2019), described one circular hotspot combined with an elongated hotspot, both in the same hemisphere, suggesting the presence of higher-order multipole moments. However, with additional data and new methodologies, Vinciguerra et al. (2024) carried out a follow-up study, and concluded that higher-order multipoles may not be necessary, but a dipole with nontrivial polar cap temperature distribution might be sufficient to explain the data. Whether higher-order multipoles are preferred in the inference method described here for this source is thus of great interest.
@@ 4.1 Inference methodology
We carry out Bayesian inference using the Markov Chain Monte Carlo (MCMC) method to find the best-fitting parameters for a shifted dipole model that can best reproduce the observed X-ray pulse profiles of the two target sources. To simplify this problem, we fix the mass and radius of these two sources to the maximum likelihood values reported in recent literature. For J0030, these values are $M=1.385M_{\odot}$ and $R=11.70\,\mathrm{km}$, as suggested by the maximum likelihood value from ST-PDT case by Vinciguerra et al. (2024). For J0437, the values are $M=1.418M_{\odot}$ and $R=11.36\,\mathrm{km}$, which corresponded to the result reported by Choudhury et al. (2024). We summarize these fixed model input value in Table 1 for J0030 and J0437 separately. This approach effectively fixes the space-time properties of the pulsar, allowing us to focus solely on the influence of the pulse profile introduced by the hotspot configuration, greatly reducing the parameter space and the amount of computational power required. A comprehensive inference that jointly infers the whole space-time properties and hotspot information will be reserved for future study.
The free parameters need to be explored include the the observer inclination $I$, magnetic inclination angle $\iota$, and the $x$–, $y$–, $z$– direction shifts in the magnetic field: $x_{0}$, $y_{0}$, and $z_{0}$. Since our model predicts only the temperature of the hotspots, we must also account for the temperature of the star’s surface outside the hotspot regions. We assume this surface temperature, $T_{s}$, to be uniform across the whole star surface. This parameter is also free and needs to be explored by our inference.
@@
The prior settings for all free parameters are summarized in Table 2. All six free parameters are set with a uniform prior (denoted as $\mathcal{U}$) to reflect the fact that we do not have a clear preference for any specific values. The $x$–, $y$–, $z$– direction shifts have a potential constraint: the total shift of the dipole field should be smaller than the star’s radius. The likelihood of the simulated light curve compared to the real data is expressed as a mean square difference error, reflecting how closely the simulated pulse profile matches the observed data. By combining the prior and likelihood, we can extract the posterior distribution and explore the allowed parameter space.
All the computations of X-ray light curve are based on the X-PSI package (Riley et al., 2023), incorporating a newly developed global temperature map module. This module theoretically handles arbitrary shapes of hotspots and temperature distributions, taking into account effects such as general relativistic light bending and rotational Doppler shift. A similar module has been implemented in Das et al. (2024). The MCMC fitting is performed using the emcee software (Foreman-Mackey et al., 2013).
Figure 7: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), $x$$y$$z$ direction shift and $T_{s}$ after applying constraints to the observed pulse profile of J0437. Blue is the posterior with Case 1: without return current. Red posterior is result in Case 2: with return current. The contour levels in the corner plot, going from deep to light colours, correspond to the 68%, 84% and 98.9% levels. The dashed line in the 1D corner plots represents the 68% credible interval, and the title of this plot indicates the median of the distribution as well as the range of the 68% credible interval. Here, $x$$y$$z$ shifts are measured in $\mathrm{km}$, while the temperature $T_{s}$ is in numerical units
Highest Likelihood Case 1 Case 2
observer inclination $I$ 0.64 0.81
magnetic inclination $\iota$ 1.36 1.14
$x_{0}\,(\mathrm{km})$1.21 -1.28
$y_{0}\,(\mathrm{km})$0.36 -0.42
$z_{0}\,(\mathrm{km})$1.32 -0.90
$\log_{10}T_{s}$ 1.89 2.51
Figure 8: (a) Polar cap temperature distribution for the maximum-likelihood hotspot configuration in Case 2: with return current for PSR J0437–4715. (b) The maximum-likelihood parameters produced pulse profile of J0437 results comparing with the data
@@ 4.2 Inference results: PSR J0030+0451
Figure 5 shows the J0030 full posterior distributions of the model parameters. Whether we perform the fitting without the return current (Case 1) or with return current (Case 2), most parameters converge to pronounced peaks in the posterior space. The observer inclination angle exhibits a consistent peak around $1.0\,\mathrm{rad}$ in both cases, suggesting that this geometry is robust to the presence or absence of return current effects. Interestingly, in Case 2 (with return current), the posterior for the observer inclination is slightly sharper. Conversely, the magnetic inclination angle $\iota$ does not converge to a single dominant peak, showing multiple local maxima. One noteworthy peak lies near $\pi/2$, indicating that placing both hotspots near the equatorial plane can plausibly reproduce the observed light curve. Analysis of the two-dimensional posterior parameter space defined by $I$ and $\iota$ in this corner plot reveals a significant degeneracy between these parameters. Specifically, when $I$ approaches $\pi/2$—indicating an equatorial viewing geometry—the magnetic inclination tends to peak at approximately $0.8\,\mathrm{rad}$. Conversely, as the magnetic inclination is toward $\pi/2$, an alternative peak emerges near $0.6\,\mathrm{rad}$. This behavior underscores the intrinsic degeneracy between the observer’s angle and the magnetic inclination. Moreover, the presence of multiple peaks in the magnetic inclination distribution suggests that its convergence is incomplete, implying that the constraints remain insufficiently precise. Notably, although both the observer angle and the magnetic inclination exhibit prominent features near the $90^{\circ}$ region, these peaks are not directly correlated, further emphasizing the fundamental degeneracy between these two parameters.
@@
For the $x$–, $y$–, and $z$–shifts, the posterior distributions reveal that the star’s dipole displacement is constrained to a relatively small fraction of the stellar radius. However, the $z$–direction shift remains less constrained than the other two, likely because its influence on the pulse profile (see Figure 4) is more moderate. In terms of the surface temperature outside the hotspot, $T_{s}$, both cases yield similar peak values, but the uncertainty range is larger without the return current included (Case 1) than the one with it included (Case 2). This result indicates that including the return current provides tighter constraints on the overall hotspot geometry. Although the more constrained parameter range in Case 2 suggests that including the return current generally provides tighter constraints on the overall hotspot geometry, our maximum-likilhood analysis indicates an magnetic inclination of approximately $90^{\circ}$. At this inclination, the impact of the return current on the hotspot geometry is minimal in this specific regime. Taken together, these findings suggest that the model with return current (Case 2) yields a more better-constrained set of parameters, implying that return current effects may be still essential for accurately modeling the observed emission.
@@
Notably, the highest peaks for the $x$–, $y$–, and $z$– shifts are all very close to zero, indicating that a minimal shift can explain the data quite well. It is also worth commenting on the inclination angle constraint. From gamma-ray observations (Abdo et al., 2013), the magnetic inclination angle presented here is consistent with other observation within 64% uncertainty range. However, given that we are fixing the mass, radius, this agreement is remarkable and interesting.
@@
In the top panel of Figure 6, we present the hotspot configuration corresponding to the maximum-likelihood sample in Case 2, with the detailed model parameters listed in Case 2 column of Figure 5. This configuration shows that both hotspots lie near the equatorial plane, and one appears slightly brighter than the other. The resulting structure is more complex than the configuration inferred by Vinciguerra et al. (2024), which did not include the temperature gradient as incorporated in our approach. This increased complexity leads to a larger magnetic inclination than the previous result, which is approximately $71.7^{\circ}$ inferred from Vinciguerra et al. (2024) ST+PDT maximum likelihood sample.
@@
This minimal model demonstrates that Case 1 and Case 2 produce similar maximum-likelihood results with comparable residuals. As the magnetic inclination approaches $90^{\circ}$, incorporating a return current produces an almost identical hotspot configuration, which is consistent with our expectations. Overall, the pulse profile of J0030 can potentially be understood within the framework of this off-center dipole field with a good fit, this result also match with the analysis in Vinciguerra et al. (2024). A comprehensive analysis, including a detailed treatment of the joint analysis of the energy spectrum and the background, will be pursued in future studies.
@@ 4.3 Inference results: PSR J0437–4715
For the magnetic inclination angle $\iota$, similar to the J0030 case, the posterior distribution for Case 2 exhibits a pronounced peak near $\pi/2$ along with additional multi-peak features that extend over a broad range, suggesting that this parameter remains poorly constrained by the current observations. The joint two-dimensional posterior distribution for the observer inclination $I$ and magnetic inclination $\iota$ reveals two distinct peaks. One solution shows the magnetic inclination approaching $\pi/2$ with an observer inclination of approximately $0.7\,\mathrm{rad}$, paralleling the behavior seen in the J0030 analysis. The alternative solution, however, features both the observer and magnetic inclinations peaking near $1.0\,\mathrm{rad}$, indicating a nontrivial configuration. Nonetheless, due to the incomplete convergence of the inference for these angles, it remains ambiguous which peak will ultimately dominate the final results. the posterior distributions for the dipole center shifts in the $x$–, $y$–, and $z$–directions are generally broader and the central value are larger compared to the J0030 inference result. This suggests that a larger dipole center shift is required to adequately explain the J0437 data. Furthermore, when comparing the scenarios with and without return current, the inclusion of return current necessitates a larger dipole shift to fit the data while simultaneously providing tighter constraints on the parameter space. This aligns with our findings from the J0030 analysis, indicating a preference for the model with return current due to its more constrained parameter space and better alignment of the maximum-likelihood observer inclination value with radio observations.
@@
All dipole-shift parameters in both cases have significantly decoupled from their priors and exhibit high-peak distributions. Regarding the surface temperature outside the hotspot, $T_{s}$, the posterior distributions are nearly identical between the two scenarios. However, for J0437, it is important to note that our inferred observer inclination is roughly consistent with the prior used in Choudhury et al. (2024), which is informed by radio observations (Reardon et al., 2024). However, if we were to restrict the prior to maintain this consistency, our inference would fail to adequately explain the observational data. There are two possible explanations for this discrepancy: first, the temperature map reported in Choudhury et al. (2024) includes a ring-shaped hotspot accompanying a circular hotspot, introducing distinct multipolar features that an off-center dipole model cannot fully capture. Second, Choudhury et al. (2024) assumes that the radio-observed orbital inclination corresponds directly to the observer inclination angle, based on the assumption that the millisecond pulsar spin axis has aligned with the spin axis of its binary system over long-term evolution. Our results suggest a more flexible assumption, allowing for a potential misalignment.
The resulting hotspot configuration, derived from the highest-likelihood sample, is illustrated in the top of Figure 8 for the case that includes return current. The detailed parameter setup is also shown in the top-right corner of Figure 7 for case 2. This configuration indicates that both the northern and southern hemispheres contain hotspots of similar area, with the northern hotspot having a slightly higher temperature. In addition, each hotspot exhibits an arched component, a feature arising from the inclusion of return current.
@@
In all of our analyses, the high computational cost of these inferences has prevented the posterior distributions from fully converging or being rigorously tested across multiple sampler configurations. The main objective of this work is to propose a new physics-motivated off-center-dipole field modeling procedure and demonstrate its feasibility using a state-of-the-art inference pipeline. Consequently, at this proof-of-concept stage, the current inference precision is considered acceptable. Future studies will undertake more advanced, large-scale inference runs that relax the currently fixed mass–radius inputs, in favor of treating the complete set of spacetime-related properties as free parameters to derive comprehensive inference once this physics-motivated framework is fully refined.
@@ 5 Discussion and Conlusion
For PSR J0030+0451, our results suggest that the inferred hotspot configuration within the context of an off-center dipole field can yield pulse profiles largely consistent with the observation data. The temperature gradient observed in our model introduces a larger magnetic inclination than that inferred from Fermi gamma-ray data (Abdo et al., 2013; Chen et al., 2020), indicating that higher-order multipole moments might still be necessary to fully account for the star’s multi-wavelength emission.
As for PSR J0437–4715, the inferred hotspot configuration is notably asymmetric, with one hotspot being significantly larger and fainter than the other. This configuration aligns with the observation data, as only one single peak is observed per rotation period. In this particular case, we see that the model with volumetric return current fits the data significantly better, indicating that it may be non-negligible at least in some scenarios.
Our results demonstrate that the off-center dipole field model, despite its simplicity, is already capable of reproducing the observed pulse profiles of both sources with a considerable degree of accuracy. However, for both sources studied in this paper, there are features that are not fully accounted by this simple model. For example, our fitted magnetic inclination for J0030 is somewhat too large, which is in tension with its gamma-ray pulse profile. Our fitted observer inclination for J0437 my not be completely consistent with the very well constrained geometry inferred from the orbital motion. Both cases likely can be further improved with a more rigorous fitting procedure that takes into account the energy-dependent pulse profile. But these inconsistencies may also point to the need for multipole magnetic configurations as those suggested by  Bilous et al. (2019),  Chen et al. (2020) and  Kalapotharakos et al. (2021).

### v3

@@ 4 Real case study: PSR J0030+0451 and PSR J0437–4715 hotspot map construction
In this case study, we focus on two sources: PSR J0030+0451 (J0030) and PSR J0437–4715 (J0437). PSR J0030+0451 was the first millisecond pulsar that NICER source for which detailed pulse profile modeling was performed. The original report on the hotspot configuration of this star, by Riley et al. (2019) and Miller et al. (2019), described one circular hotspot combined with an elongated hotspot, both in the same hemisphere, suggesting the presence of higher-order multipole moments. However, with new background constraint from XMM-Newton data, Vinciguerra et al. (2024) carried out a follow-up study, and concluded that higher-order multipoles may not be necessary, but a dipole with nontrivial polar cap temperature distribution might be sufficient to explain the data. Whether higher-order multipoles are preferred in the inference method described here for this source is thus of great interest.
@@ 4.1 Inference methodology
We carry out Bayesian inference using the Markov Chain Monte Carlo (MCMC) method to find the best-fitting parameters for a shifted dipole model that can best reproduce the observed X-ray pulse profiles of the two target sources. To simplify this problem, we fix the mass and radius of these two sources to the maximum likelihood values reported in recent literature. For J0030, these values are $M=1.385M_{\odot}$ and $R=11.70\,\mathrm{km}$, as suggested by the maximum likelihood value from ST-PDT case by Vinciguerra et al. (2024). For J0437, the values are $M=1.418M_{\odot}$ and $R=11.36\,\mathrm{km}$, which corresponded to the result reported by Choudhury et al. (2024). We summarize these fixed model input value in Table 1 for J0030 and J0437 separately. This approach effectively fixes the space-time properties of the pulsar, allowing us to focus solely on the influence of the pulse profile introduced by the hotspot configuration, greatly reducing the parameter space and the amount of computational power required. The MCMC chains presented below are exploratory just sufficient to demonstrate that the physics-motivated / shifted-dipole model presented here can be sampled with a standard Bayesian pipeline. We have not yet established statistical convergence across alternative sampling schemes, even though the chains display qualitative convergence indicators, and Figures 7 still show minimal visible sampling artifacts despite cosmetic changes (wider axes and KDE contours) adopted to minimize distraction. A full-scale analysis with longer chains, multiple sampler configurations, and the mass–radius parameters set free is deferred to future work.
The free parameters need to be explored include the the observer inclination $I$, magnetic inclination angle $\iota$, and the xx–, yy–, zz– direction shifts in the magnetic field: $x_{0}$, $y_{0}$, and $z_{0}$. Since our model predicts only the temperature of the hotspots, we must also account for the temperature of the star’s surface outside the hotspot regions. We assume this surface temperature, $T_{s}$, to be uniform and finite across the whole star surface. This parameter is also free and needs to be explored by our inference.
@@
The prior settings for all free parameters are summarized in Table 2. All six free parameters are set with a uniform prior (denoted as $\mathcal{U}$) to reflect the fact that we do not have a clear preference for any specific values. The xx–, yy–, zz– direction shifts have a potential constraint: the total shift of the dipole field should be smaller than the star’s radius. The likelihood of the simulated light curve compared to the real data is expressed as a mean square difference error, reflecting how closely the simulated pulse profile matches the observed data. By combining the prior and likelihood, we can extract the posterior distribution and explore the allowed parameter space.
All the computations of X-ray light curve are based on the X-PSI package (Riley et al., 2023), incorporating a newly developed global temperature map module. This module theoretically handles arbitrary shapes of hotspots and temperature distributions, taking into account effects such as general relativistic light bending and rotational Doppler shift. Here we used blackbody as the emission model without including realistic atmosphere model to reduce the computation time. A similar module has been implemented in Das et al. (2024). The MCMC fitting is performed using the emcee software (Foreman-Mackey et al., 2013).
Figure 7: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), xxyyzz direction shift and $T_{s}$ after applying constraints to the observed pulse profile of J0437. Blue is the posterior with Case 1: without return current. Red posterior is result in Case 2: with return current. The contour levels in the corner plot, going from deep to light colours, correspond to the 68%, 84% and 98.9% levels. The dashed line in the 1D corner plots represents the 68% credible interval, and the title of this plot indicates the median of the distribution as well as the range of the 68% credible interval. Here, xxyyzz shifts are measured in $\mathrm{km}$, while the temperature $T_{s}$ is in numerical units
Maximum Likelihood Case 1 Case 2
observer inclination II 0.40 0.59
magnetic inclination $\iota$ 1.22 1.09
$x_{0}\,(\mathrm{km})$ 1.90 -1.85
$y_{0}\,(\mathrm{km})$ -0.67 -2.07
$z_{0}\,(\mathrm{km})$ -0.27 -0.29
$\log_{10}T_{s}$ 2.33 2.53
Figure 8: (a) Polar cap temperature distribution for the maximum-likelihood hotspot configuration in Case 2: with return current for PSR J0437–4715. (b) The maximum-likelihood parameters produced pulse profile of J0437 results comparing with the bolometric flux data in NICER band. A companion animated figure showing the evolution of this hotspot configuration as the off‑center distance increases toward the best‑fit state is available online; an archived copy is hosted on Zenodo: https://doi.org/10.5281/zenodo.16548609 (catalog doi:10.5281/zenodo.16548609).
@@ 4.2 Inference results: PSR J0030+0451
Figure 5 shows the J0030 full posterior distributions of the model parameters. Whether we perform the fitting without the return current (Case 1) or with return current (Case 2), most parameters converge to pronounced peaks in the posterior space. The observer inclination angle exhibits a consistent peak around $1.0\,\mathrm{rad}$ in both cases, suggesting that this geometry is robust to the presence or absence of return current effects. The posterior distribution of the magnetic inclination angle, $\iota$, converges to a single dominant peak near $1.0$, indicating satisfactory convergence even without repeated runs under alternative sampling configurations.
@@
For the $x$-, $y$-, and $z$-shift parameters, the posterior distributions indicate that, for J0030+0451, the stellar dipole is displaced by only a small fraction of the stellar radius: all shifts are well constrained within 1 km (i.e., $10\%$ of $R_{\star}$). In terms of the surface temperature outside the hotspot, $T_{s}$, both cases yield similar peak values, but the uncertainty range is larger without the return current included (Case 1) than the one with it included (Case 2). This result indicates that including the return current provides tighter constraints on the overall star surface background temperature. Our maximum-likelihood analysis favors a magnetic inclination in which the model without a return current yields $\iota\approx 90^{\circ}$, whereas the model including a return current gives a more plausible value of $\iota\sim 56^{\circ}$. This contrast strongly supports the necessity of incorporating return-current physics when using first-principles, physics-motivated hotspot models to interpret the observations. Taken together, these findings suggest that the model with return current (Case 2) yields a more reasonable set of parameters, implying that return current effects may be essential for accurately modeling the observed emission.
@@
Notably, the highest peaks for the xx–, yy–, and zz– shifts are all very close to zero, indicating that a minimal shift can explain the data quite well. It is also worth commenting on the inclination angle constraint. From gamma-ray observations (Abdo et al., 2013), the magnetic inclination angle presented here is smaller other observation within 64% uncertainty range. However, given that we are fixing the mass, radius, this is still remarkable and interesting.
@@
In the top panel of Figure 6, we present the hotspot configuration corresponding to the maximum-likelihood sample in Case 2, with the detailed model parameters listed in Case 2 column of Figure 5. This configuration shows that both hotspots lie near the equatorial plane, and one appears significantly brighter but smaller than the other. The resulting structure is more complex than the configuration inferred by Vinciguerra et al. (2024), which did not include the temperature gradient as incorporated in our approach. This increased complexity leads to a smaller magnetic inclination than the previous result, which is approximately $71.7^{\circ}$ inferred from Vinciguerra et al. (2024) ST+PDT maximum likelihood sample.
@@
This minimal model demonstrates that Case 1 and Case 2 produce similar maximum-likelihood results with comparable residuals. Overall, the pulse profile of J0030 can potentially be understood within the framework of this off-center dipole field with a good fit, this result also match with the analysis in Vinciguerra et al. (2024). A comprehensive analysis, including a detailed treatment of the joint analysis of the energy spectrum and the background, will be pursued in future studies.
@@ 4.3 Inference results: PSR J0437–4715
For the magnetic inclination angle $\iota$, similar to the J0030 case, the posterior distribution for Case 2 exhibits a pronounced peak near $1.0$ with a wide peak, suggesting that this parameter remains less constrained by the current observations. The 2D posterior distribution features both the observer and magnetic inclinations peaking near $1.0\,\mathrm{rad}$, indicating a nontrivial configuration. The posterior distributions for the dipole center shifts in the xx–, yy–, and zz–directions are generally broader and the central value are larger compared to the J0030 inference result. This suggests that a larger dipole center shift is required to adequately explain the J0437 data. Furthermore, in direct comparisons between models with and without a return current, the inclusion of a return current can help the dipole-shift posterior contract to a predominant peak and simultaneously yields tighter constraints across the parameter space, whereas, in the case without a return current, the posterior is far more dispersed and exhibits poor convergence, further highlighting the importance of including a return current. Collectively, these results indicate improved identifiability of the parameters when return-current physics is incorporated. This aligns with our findings from the J0030 analysis, indicating a preference for the model with return current due to its more constrained parameter space and better alignment of the maximum-likelihood observer inclination value with radio observations.
@@
All dipole-shift parameters in both models are strongly decoupled from their priors and exhibit sharply peaked posteriors. For the surface temperature outside the hotspot, $T_{s}$, the posterior distributions are nearly identical across the two scenarios, however, in the model without a return current the distribution is noisier, with some samples extending to very large $T_{s}$ and a small secondary peak near $T_{s}\sim 4$. However, for J0437, it is important to note that our inferred observer inclination is roughly consistent with the prior used in Choudhury et al. (2024), which is informed by radio observations (Reardon et al., 2024). However, if we were to restrict the prior to maintain this consistency, our inference would fail to adequately explain the observational data. There are two possible explanations for this discrepancy: first, the temperature map reported in Choudhury et al. (2024) includes a ring-shaped hotspot accompanying a circular hotspot, introducing distinct multipolar features that an off-center dipole model cannot fully capture. Second, Choudhury et al. (2024) assumes that the radio-observed orbital inclination corresponds directly to the observer inclination angle, based on the assumption that the millisecond pulsar spin axis has aligned with the spin axis of its binary system over long-term evolution. Our results suggest a more flexible assumption, allowing for a potential misalignment.
The resulting hotspot configuration, derived from the highest-likelihood sample, is illustrated in the top of Figure 8 for the case that includes return current. The detailed parameter setup is also shown in the top-right corner of Figure 7 for case 2. This configuration indicates that the northern hotspot exhibits a ring-shaped bright pattern, whereas the southern hotspot spans a larger area but is substantially fainter. Moreover, an additional asymmetry appears as a small emission region in the southeastern quadrant. This morphology is particularly intriguing, since the best-fitting result of Choudhury et al. (2024) likewise favors a configuration that is qualitatively similar and features an asymmetric hotspot pattern. To better demonstrate how an off-centered dipole field yields the observed asymmetric hotspot geometry, we provide a companion video that visualizes the evolution of the best-fitting hotspot configuration as the off-center distance is increased to its optimal value. The video is available on Zenodo: https://doi.org/10.5281/zenodo.16548609 (catalog doi:10.5281/zenodo.16548609). In addition, each hotspot exhibits an important arched component, a feature only arising from the inclusion of return current.
@@
In all of our analyses, the high computational cost of these inferences has prevented being rigorously tested across multiple sampler configurations. The main objective of this work is to propose a new physics-motivated off-center-dipole field modeling procedure and demonstrate its feasibility using a state-of-the-art inference pipeline. Consequently, at this proof‑of‑concept stage, the parameter estimates should be regarded as illustrative rather than definitive: the chains achieve a plausibly reasonable precision and display apparent convergence, yet a rigorous, multi sampling setup convergence demonstration is deferred to future work. Future studies will undertake more advanced, large-scale inference runs that relax the currently fixed mass–radius inputs, in favor of treating the complete set of spacetime-related properties as free parameters to derive comprehensive inference once this physics-motivated framework is fully refined.
@@ 5 Discussion and Conlusion
For PSR J0030+0451, our results suggest that the inferred hotspot configuration within the context of an off-center dipole field can yield pulse profiles largely consistent with the observation data. The temperature gradient observed in our model introduces a smaller magnetic inclination than that inferred from Fermi gamma-ray data (Abdo et al., 2013; Chen et al., 2020), indicating that higher-order multipole moments might still be necessary to fully account for the star’s multi-wavelength emission.
As for PSR J0437–4715, the inferred hotspot configuration is notably asymmetric, with one hotspot exhibiting an interesting asymmetric structure that includes a small component located in the southeastern quadrant of the hotspot. This configuration surprisingly aligns with the best-fit hotspot configuration reported by Choudhury et al. (2024). In this particular case, we see that the model with volumetric return current fits the data significantly better, indicating that it may be non-negligible at least in some scenarios.
Our results demonstrate that the off-center dipole field model, despite its simplicity, is already capable of reproducing the observed pulse profiles of both sources with a considerable degree of accuracy. However, for both sources studied in this paper, there are features that are not fully accounted by this simple model. For example, our fitted magnetic inclination for J0030 is somewhat small, which is in tension with its gamma-ray pulse profile. Our fitted observer inclination for J0437 not be completely consistent with the very well constrained geometry inferred from the orbital motion. Both cases likely can be further improved with a more rigorous fitting procedure that takes into account the energy-dependent pulse profile. But these inconsistencies may also point to the need for multipole magnetic configurations as those suggested by  Bilous et al. (2019),  Chen et al. (2020) and  Kalapotharakos et al. (2021).